OpenStudy (anonymous):

domain of 2x^2-3x-9 domain of 2x^2-3x-9 @Mathematics

OpenStudy (anonymous):

all real numbers

OpenStudy (anonymous):

like?

OpenStudy (anonymous):

solve it out 1st? pls

jimthompson5910 (jim_thompson5910):

The domain of any polynomial is the set of all real numbers (since you can plug in any real number for x and get a result). There's not much more to it than that.

OpenStudy (hoblos):

\[\mathbb{R}\]

OpenStudy (anonymous):

all real numbers, anything u plug into x will give u a real number

OpenStudy (anonymous):

is it the same as -infinity to infinity?

jimthompson5910 (jim_thompson5910):

yes, this is representing the answer in interval notation.

OpenStudy (hoblos):

yup

jhonyy9 (jhonyy9):

x_1,_2=(3+/- sqrt(9+72))/4 =(3+/- 9)/4 = -6/4 = -3/2 and 3 - so those are zeros of this functions